OpenStudy (anonymous):

Please help!! will give fan and medal(: Which of the following could be the sides of a right triangle? A. 5, 6, and 8 B. 12, 13, and 18 C. 15, 16, and 22 D. 20, 21, and 29

OpenStudy (cwrw238):

check which fit the pythagoras formula a^2 = b^2 + c^2 where a = hypotenuse (longest side)

OpenStudy (anonymous):

can you show me? please

OpenStudy (anonymous):

@cwrw238